Posters.com is a small Internet retailer of high-quality posters. The company has $840,000 in operating assets and fixed expenses of $162,000 per year. With this level of operating assets and fixed expenses, the company can support sales of up to $5,000,000 per year. The company’s contribution margin ratio is 10%, which means that an additional dollar of sales results in additional contribution margin, and net operating income, of 10 cents. Required: 1. Complete the following table showing the relation between sales and return on investment (ROI). 2. What happens to the company’s return on investment (ROI) as sales increase?

Complete the following table showing the relation between sales and return on investment (ROI). (Round your percentage answers to 2 decimal places.)

Sales Net Operating Average Operating ROI
Income Assets %
$4,500,000 $288,000 $840,000 %
$4,600,000 $840,000 %
$4,700,000 $840,000 %
$4,800,000 $840,000 %
$4,900,000 $840,000 %
$5,000,000 $840,000 %
Answer #1

Complete the following table showing the relation between sales and return on investment (ROI). (Round your percentage answers to 2 decimal places.)

Sales Net Operating Average Operating ROI
Income Assets %
$4,500,000 $288,000 $840,000 288000/840000 = 34.29 %
$4,600,000 298000 $840,000 298000/840000 = 35.48 %
$4,700,000 308000 $840,000 308000/840000 = 36.67 %
$4,800,000 318000 $840,000 318000/840000 = 37.86 %
$4,900,000 328000 $840,000 328000/840000 = 39.05 %
$5,000,000 338000 $840,000 338000/840000 = 40.24 %